# OODTE: A Differential Testing Engine for the ONNX Optimizer

### v1

@@ IV Related Work
Overall, limited work has been done in the direction of testing AI compiler optimizations. Choi et al. [14] have conducted an empirical study towards optimizer comparisons, however they do not consider graph-based optimizations in the post-training phase such as the ONNX Optimizer) as they focus on stochastic optimizers focusing primarily on the model training process, such as Adam [29]. DiffChaser [73] explores the application of differential testing across optimized and unoptimized DNN model variants. Louloudakis et al. propose DeltaNN [41], a utility to explore model robustness across different configurations - including but not limited to compiler optimizations. Xiao et al. [72] propose a metamorphic testing methodology for AI compiler bug detection, including issues related to optimizations.
@@ V Experiment Setup
For our experiment set, we sourced $130$ model instances from the ONNX Model Hub. In particular, we selected $93$ models utilized for classification tasks, $30$ models used for object detection and semantic segmentation tasks, and $7$ transformer models [54] utilized for text generation tasks, such as question and answering, (BERT-SQuAD [18]), sentiment analysis (RoBERTa [40]), text summarization (T5 [51]) and generic-purpose text generation (GPT-2 [50]). We utilize models of a variety of opsets based on hub availability, ranging from $7$ (the lowest compatible with the latest ONNX Runtime version by the time of this publication - ONNX Runtime v1.21.11.21.11.21.11.21.1) up to $12$ (the highest provided from ONNX Model Hub via their API). However, our approach is generalizable and not restricted to one hub or limited by opset versioning. We selected it because it consists of a standard set of models, refined and prepared for wide use by developers and researchers. The same applies for the models we selected: we focused on models that are mature and strong representatives of their categories (e.g., MobileNet [24], EfficientNet-Lite4 [65] DenseNet121 [25], InceptionV2 [63] RetinaNet [36], Mask/Faster R-CNN R-50-FPN [21, 56] SSD [39], RoBERTa [40] and GPT-2 [50]), but our methodology can apply to any type of ONNX model. In addition, we utilized text generation models that generated results in a direct, one-off manner (such as RoBERTa), but also in a token retrofitting manner (such as GPT-2). In the latter case, we considered selection of the first, second, and third token prediction, so that our results are as comprehensive as possible.
In terms of computational resources, we ran our models on mid-range systems with Intel i5 and i7 CPUs, reflecting the typical hardware used by developers during model development and fine-tuning. Our method is generalizable to larger models, limited only by ONNX model size constraints, and can be extended to high-end systems with a trivial extension for hardware acceleration (using the ONNX Runtime API [3], or an AI compiler such as Apache TVM [12]).
We selected datasets appropriate to each model task. For classification and object detection, we used the ILSVRC2017 [57] validation dataset, containing $5500$ images, commonly used for benchmarking image recognition models. For text generation, we used SQuAD-v2 [52], which includes $11873$ questions and contexts, to evaluate Q&A tasks and general-purpose transformer behavior. Additionally, for summarization and sentiment analysis, we used the IMDB dataset [44, 66], containing $25000$ entries.
@@
    For sentiment analysis, we compared each output value consisting of ${0,1}$ across the original and the optimized variants, registering any differences observed.
Finally, we evaluated all existing ONNX Optimizer passes. However, OODTE can also be applied to new or custom passes. Moreover, our methodology can be extended to other optimizers or model optimization APIs (e.g., Apache TVM), provided that the API supports applying passes both individually or as a bundle.
@@ VI Results
Following the execution of the full set of experiments, OODTE successfully identified 15 bugs. Of these, 14 were entirely new (not previously reported on the ONNX Optimizer issue tracker), while one was a previously reported issue that lacked sufficient detail, including model information and reproducibility steps. We provided comprehensive information and full reproducibility details for this issue.
The issues were associated with $9$ of the $47$ passes in the optimizer. However, not all passes are of equal importance; by default, the optimizer utilizes passes related to fuse and eliminate operations. Therefore, $8$ of the discovered issues are deemed more critical, as they are part of the optimizer’s core bundle (default) strategy. Of these, $6$ are directly linked to specific passes (e.g., fuse_bn_to_conv, which fuses batch normalization nodes with convolutions), while the remaining issues were general and related to the optimizer’s overall usage. We present the impact of these faults both on the optimization process and on the correctness of the optimized models.
TABLE I: Issues detected using OODTE, along with models manifested at. Opset version indicated near each model name.
@@
1 lift_lexical_references Removes used value references. Crash EfficientNet-Lite4 (11), SSD (12)
2 split_init List index out of range Crash YOLOv3 (10, 12)
3 split_init split_predict Field is required but missing. Crash DenseNet121 (7-9), YOLOv2 (10), SSD (12), Inception v1 & v2 (7-9), Tiny-YOLOv3 (11) YOLOv3 (10, 12)
4 split_predict Required inputs are missing. Crash SSD (12), DenseNet121 (7), InceptionV2 (6-9)
5 split_predict Bad spec for node. Crash YOLOv3 (10), DenseNet121 (7) InceptionV2 (6-9)
6 split_predict Input element number mismatch. Crash FCN ResNet-50-int8 (12)
7 eliminate_shape_gather Indices value not equal to node dims size Crash Faster R-CNN R-50-FPN (10)
8 All Passes Models present slightly different results. Diff. See Table II
9 rewrite_input_dtype Output name used multiple times. Crash GPT-2 (10)
10 eliminate_ consecutive_ idempotent_ops Incompatible Shape Error. Crash (Run) Mask R-CNN/Faster R-CNN: R-50-FPN-int8 (12), R-50-FPN-qdq (12), R-50-FPN-fp32 (12)
11 eliminate_ duplicate_ initializer Erase output assertion failed. Crash Faster R-CNN R-50-FPN-qdq (12)
12 fuse_bn_into_conv Duplicate initializer introduced. Crash (Run) RetinaNet (9)
13 eliminate_shape_gather eliminate_nop_reshape Added initializers not used by the model. Warning Faster R-CNN R-50-FPN-int8 (12)
14 All Passes Invalid input scale. Crash VGG-16-BN (7), VGG-19-BN (7), ResNet101_DUC_HDC (7)
15 fuse_bn_into_conv Bad graph structure Diff. See Table II
@@ VI-A Optimizer Crashes
Out of the $15$ issues, $12$ were related to model crashes, $2$ caused model accuracy deviations and $1$ obfuscated the model with redundant initializers, leading to a runtime warning. In particular, we observed that those crashes concerned $8$ out of the $47$ passes ($17$%), and occurred for specific models, with object detection models being more prone to errors. In particular, R-CNN models, such as Faster R-CNN R-50-FPN [56], RetinaNet [36], VGG-16, VGG-19 [60], as well as YOLOv3 [55], presented crashes by using the optimizer, in a range of opset instances and in passes such related to the modification of gather nodes, as well as the elimination of idempotent operations. In total, we observed $31$ model instances that either the optimizer crashed upon processing them, or the optimized model was problematic, leading into a runtime crash. However, some secondary passes are known to be unstable (e.g., split_init and split_predict describe in their documentation that they not strictly comply with the ONNX format), and therefore we do not focus on them as sources of faults in the optimization process. As a result, from those, $12$ model instances ($9.2$%) while using the primary (fuse and eliminate) ONNX Optimizer passes and are of higer concern.
Overall, classification models were affected towards optimizer crashes to a much lesser extent. However, they were considerably affected in terms of correctness by a subset of faults that did not lead to a crash - which we present below. From text generation models, one issue was related to the crash of the optimizer, and it concerned the rewriting of the input type. However, other text generation models did not present any crash behaviors while optimized using the ONNX optimizer.
VI-B Deviations In Optimized Model Accuracy
Out of the $130$ model variants tested, we observed unexpected accuracy differences between the original and optimized models. This was an unexpected outcome, since ONNX Optimizer transformations are graph-based and do not alter numeric precision, meaning no accuracy deviations were anticipated. Specifically, of the $93$ classification models, $28$ showed accuracy differences, while $5$ of $30$ object detection models had discrepancies. Only one transformer instance (GPT-2) exhibited negligible differences ($<0.001\%$), with all others showing identical results. The differences primarily affected the top-5 and top-10 predictions, with no changes in the top-1 predictions, except for the SSD model, which showed a $2.9\%$ difference in top-1 prediction. Although top-1 accuracy is often prioritized, it is important to recognize that changes in lower-ranked predictions can also indicate issues. Our methodology identified these inconsistencies, which have been reported to the optimizer developers. For instance, ResNet50-caffe2 showed a $12.9\%$ difference in top-10 labels and $2\%$ in top-5 labels, while ShuffleNet presented $0.5\%$ and $1\%$ differences in top-5 and top-10 predictions, respectively.
Figure 2: Illustrative example of problematic label classification and object detection between original (blue) and optimized (orange) model variants, with the bounding box being affected in both cases, with the same (left) and different (right) object label prediction, along with their affected certainty scores.
Figure 3: Model graph malformation caused by the fuse_bn_to_conv pass in EfficientNet-Lite-4.
@@ VI-B Deviations In Optimized Model Accuracy
TABLE II: Observed differences across classification predictions for classification (CL) and object detection (OD) models.
Model Type Opset(s) Top 1 Top 5 Top 10 Issue
AlexNet [30] CL 7-9 0% 0% 0.02% Format Version Change
DenseNet121 [25] CL 7-9 0% 0% 0.02% Format Version Change
EfficientNet-Lite4 [65] CL 11 0% 0.04% 1% Graph Malformation
GoogleNet [64] CL 7-9 0% 0% 0.04% Format Version Change
Inception-1 [63] CL 7-9 0% 0% 0.25% Format Version Change
Inception-2 [63] CL 7-9 0% 0.02% 0.05% Format Version Change
MobileNet v2 [24, 58] CL 7-9 0% 0% 0.02% Format Version Change
ResNet18 [22] CL 7 0% 0.02% 0.02% Format Version Change
ResNet34-v2 [23] CL 7 0% 0.02% 0.02% Format Version Change
ResNet50-caffe2 [22] CL 7-9 0% 2% 13% Format Version Change
ShuffleNet-v1 [75] CL 7-9 0% 0.7% 1% Format Version Change
SqueezeNet 1.0 [26] CL 7-9 0% 0% 0.02% Format Version Change
SSD [39] OD 12 2.9% 4.3% 6% Graph Malformation
Following runs with OODTE, we concluded that $2$ primary issues were related with accuracy deviations across models:
@@
  • The optimizer automatically upgrades ONNX formats of $v3$ to $v4$, without any warning or indication. This affects models of opsets $7$, $8$, and $9$, which is an outdated opset, but still backwards compatible to the current ONNX Runtime version (v1.21.1), well-represented in the ONNX Model Hub, with $41\%$ of the models in it consisting of any of these opset versions. Of course, not all models within this opset range are affected, yet we found a considerable amount of models presenting differences across their original and optimized variants. We have reported this issue to the optimizer developers and they have acknowledged this behavior for further inspection.
  • The pass fuse_bn_to_conv essentially fuses BatchNormalization and Convolution operators within the model. We observed that this pass was found problematic in a number of cases, resulting in both a crashing bug and an issue where redundant nodes are introduced to the model. In the latter case, the pass attempts to fuse BatchNormalization and Conv nodes, but does so in an incorrect manner, essentially breaking both node properties into chunk nodes. The result can be visually observed in Figure 3. We found that a number of models were producing different accuracies when optimized using this pass. This issue was previously encountered by another user of the optimizer, however no details were given regarding how this issue could be reproduced or what models were affected by it. We raised a comprehensive issue to the ONNX Optimizer developers, providing all the necessary information and details related to its behavior.
Table II showcases the accuracy differences detected in top-K labels for classification and object detection moels, indicating which of the above issues is responsible for this behavior. As an example, Figure 2 illustrates how misclassification and bounding box issues manifest in an input image, with relation to the original (blue) and the optimized (orange) model.
@@
(a) Object Detection Scores with IoU threshold of 50%.
@@
(b) Object Detection Scores with IoU threshold of 75%.
@@
(c) Object Detection Scores with IoU threshold of 90%.
VI-C Optimizer Warnings
We observed two issues related to the obfuscation of the optimized models with unused initializer entries, associated with the eliminate_shape_gather and eliminate_nop_reshape passes. When executed with ONNX Runtime, these entries were ignored, triggering a warning in the format: Removing initializer ’INITIALIZER_ID’. It is not used by any node and should be removed from the model. Upon manual inspection, we found no significant differences in model size. However, this issue warrants further investigation for potential side effects. We have reported it to the optimizer developers for additional review.

### v2

@@ IV Related Work
Overall, limited work has been done in the direction of testing AI compiler optimizations. Choi et al. [14] have conducted an empirical study towards optimizer comparisons, however they do not consider graph-based optimizations in the post-training phase such as the ONNX Optimizer) as they focus on stochastic optimizers focusing primarily on the model training process, such as Adam [29]. DiffChaser [72] explores the application of differential testing across optimized and unoptimized DNN model versions. Louloudakis et al. propose DeltaNN [41], a utility to explore model robustness across different configurations - including but not limited to compiler optimizations. Xiao et al. [71] propose a metamorphic testing methodology for AI compiler bug detection, including issues related to optimizations.
@@ V Experiment Setup
V-1 The Experiment Process
We evaluated all ONNX Optimizer passes (as detailed in Section II-C). For each original model under test, we first applied the default setting of the ONNX Optimizer, which applies all fusion and elimination passes at once, generating the optimized model. We then compared the outputs of the original and optimized models according to the model type, as described in Section III-D. For optimized models presenting output differences, we performed optimization using only a single optimization pass, resulting in $47$ such single-pass optimized models. To identify which pass causes the problematic behavior, we compared the outputs of each single-pass model against those of the original model. It is worth noting that OODTE is compatible with new or custom passes as well. Furthermore, our methodology generalizes to other optimization frameworks (e.g., Apache TVM), provided they support both bundled and individual pass applications.
V-2 DNN Models
For our experiment set, we sourced $130$ model instances from the ONNX Model Hub [7]. In particular, we selected $93$ models utilized for classification tasks (e.g., MobileNet [24], EfficientNet-Lite4 [63] DenseNet121 [25], and InceptionV2 [61]), $30$ models used for object detection and semantic segmentation tasks (e.g., RetinaNet [36], Mask/Faster R-CNN R-50-FPN [21, 54], and SSD [39]), and $7$ transformer models [66] utilized for text comprehension and generation tasks, such as question answering (Q&A), (BERT-SQuAD [18]), sentiment analysis (RoBERTa [40]), text summarization (T5 [50]) and generic-purpose text generation (GPT-2 [49]), with $4$ models focusing on Q&A, while one model is utilized for each of the other text comprehension and generation tasks. We utilized models of a variety of opsets based on hub availability, ranging from $7$ (the lowest compatible with the latest ONNX Runtime version by the time of this publication - v1.21.11.21.11.21.11.21.1) up to $12$ (the highest provided from ONNX Model Hub via their API). Our methodology, however, is applicable to any ONNX model type and is not constrained by opset versions or tied to a specific model repository.
V-3 Computational Resources
We ran our models on mid-range systems with Intel i5 and i7 CPUs, reflecting the typical hardware used by developers during model development and fine-tuning. Our method is generalizable to larger models, limited only by ONNX model size constraints, and can be extended to high-end systems with a trivial extension for hardware acceleration (using the ONNX Runtime API [2], or an AI compiler such as Apache TVM [12]).
V-4 Datasets
We selected datasets appropriate to each model task. For classification and object detection, we used the ILSVRC2017 [55] validation dataset, containing $5500$ images, commonly used for benchmarking image recognition models. For text comprehension and generation, we used SQuAD-v2 [51], which includes $11873$ questions and contexts, to evaluate Q&A tasks and general-purpose transformer behavior. Additionally, for summarization and sentiment analysis, we used the IMDB dataset [44, 64], containing $25000$ entries.
@@ V-4 Datasets
    For sentiment analysis models, we compared each output value consisting of ${0,1}$ across the original and the optimized versions, registering any differences observed.
@@ VI Results
Following the execution of the full set of experiments11footnotemark: 1, OODTE successfully identified $15$ bugs, reported across $11$ separate GitHub issues [9]. A complete list of these issues, along with the associated models, is shown in Table I. Of the 15 bugs identified, $14$ were entirely new (not previously reported on the ONNX Optimizer issue tracker), while one was a previously reported issue that lacked sufficient detail, including model information and reproducibility steps. We provided comprehensive information and full reproducibility details for this issue.
Within the ONNX optimizer, the bugs identified were primarily associated with $9$ of the $47$ passes. However, not all passes are equally important. By default, the optimizer prioritizes passes related to fuse and eliminate operations. As a result, $8$ of the $11$ reported GitHub issues are considered more critical since they belong to the optimizer’s core (default) strategy. Among these $8$ issues, $6$ are linked to specific passes (e.g., fuse_bn_to_conv, which fuses batch normalization nodes with convolutions), while the remaining 2 issues were more general and related to the optimizer’s overall usage. We analyze and present the impact of these faults both on the optimization process and on the correctness of the optimized models in the following Sections.
TABLE I: Overview of the issues identified using OODTE, detailing the responsible optimization pass, the detected error, its type (crash or accuracy), and the models in which the error was observed, along with their associated opsets (op.).
@@
1 lift_lexical_references Removes used value references. Crash EfficientNet-Lite4 (op. 11), SSD (op. 12)
2 split_init List index out of range. Crash YOLOv3 (op. 10, 12)
3 split_init split_predict Field is required but missing. Crash DenseNet121 (op. 7-9), YOLOv2 (op. 10), SSD (op. 12), Tiny-YOLOv3 (op. 11) YOLOv3 (op. 10, 12)
4 split_predict Required inputs are missing. Crash SSD (op. 12), DenseNet121 (op. 7), InceptionV2 (op. 6-9)
5 split_predict Bad spec for node. Crash YOLOv3 (op. 10), DenseNet121 (op. 7) InceptionV2 (op. 6-9)
6 split_predict Input element number mismatch. Crash FCN ResNet-50-int8 (op. 12)
7 eliminate_shape_gather Indices value not equal to node dims size Crash Faster R-CNN: R-50-FPN (op. 10)
8 All Passes Implicit ONNX format version change. Accuracy Models of cases $1$-$11$ in Table II
9 rewrite_input_dtype Output name used multiple times. Crash GPT-2 (op. 10)
10 eliminate_ consecutive_ idempotent_ops Incompatible shape error. Crash (Run) Mask R-CNN/Faster R-CNN: R-50-FPN-int8 (12), R-50-FPN-qdq (op. 12), R-50-FPN-fp32 (op. 12)
11 eliminate_ duplicate_ initializer Erase output assertion failed. Crash Faster R-CNN: R-50-FPN-qdq (12)
12 fuse_bn_into_conv Duplicate initializer introduced. Crash (Run) RetinaNet (op. 9)
13 eliminate_shape_gather eliminate_nop_reshape Added initializers not used by the model. Warning Faster R-CNN: R-50-FPN-int8 (op. 12)
14 All Passes Invalid input scale. Crash VGG-16-BN (op. 7), VGG-19-BN (op. 7), ResNet101_DUC_HDC (op. 7)
15 fuse_bn_into_conv Bad graph structure. Accuracy Efficient-Lite4 (op. 11), SSD (op. 12)
@@ VI-A Optimizer Crashes
Out of the $15$ detected bugs, $12$ bugs were related to model crashes, $2$ bugs resulted in output differences between the original and the optimized model, and $1$ bug introduced redundant initializers in the model that produced runtime warnings. In particular, we observed crashes were caused by $8$ out of the $47$ passes ($17$%), and occurred for specific models, with object detection models being more prone to errors from these $8$ passes. In particular, R-CNN models, such as Faster R-CNN R-50-FPN [54], RetinaNet [36], VGG-16, VGG-19 [58], as well as YOLOv3 [53], presented crashes by using the optimizer, in a range of opset instances and in passes related to the modification of gather nodes, and the elimination of idempotent operations. In total, we observed $31$ model instances where either the optimizer crashed upon processing them, or the optimized model was problematic, leading to a runtime crash. Some secondary passes are known to be unstable (e.g., split_init and split_predict describe in their documentation that they not strictly comply with the ONNX format), and therefore we do not focus on them as sources of faults in the optimization process. Overall, $12$ out of the $31$ model instances ($9.2$%) that were associated with a crash occurred while using passes related to fuse and eliminate operations, which are part of the default bundle of optimizations in the ONNX Optimizer and are therefore of particular significance.
As shown in Table I, crash faults can be attributed to three main causes: (1) model input issues (cases #$4$ and #$6$), (2) incorrect use of initializers (cases #$12$ and #$13$), and (3) issues associated with node properties such as duplicate names, obsolete fields, and incorrect dimensions (all other cases). Overall, object detection models were more prone to optimizer crashes compared to classification models. Due to their increased structural complexity and larger size, they are more likely to exhibit errors related to node properties and initializers, which account for the vast majority of crash faults. In contrast, only one issue was observed in the most complex model category — text generation models, that was related to model input in GPT-2. However, upon further inspection, we found that the transformation passes of the ONNX Optimizer had limited impact on text generation models — the majority of optimization passes left the models unchanged, and those that did make modifications affected a small proportion of nodes with negligible effect on outputs.
@@
VI-B Output Label Differences in Optimized Classification & Object Detection Models
Out of the $130$ optimized models tested, we observed unexpected output differences between the original and optimized models. These differences manifested as (1) label differences for classification models (e.g., ResNet), (2) label and/or bounding box differences for object detection models (e.g., YOLOv3), and (3) textual output differences in token sequences or predicted spans for text generation tasks. Since the ONNX Optimizer transformations are graph-based and do not alter numeric precision, output differences are not expected. Among the $93$ classification models tested, $28$ ($30.1$%) exhibited discrepancies in output labels. Similarly, $5$ out of $30$ ($16.6$%) object detection models presented output differences. The text generation models, on the other hand, were robust with only one instance (GPT-2) exhibiting negligible differences ($<0.001\%$) while all others showing identical results across the original and the optimized model versions. The differences primarily affected the top-5 and top-10 predictions, with no changes in the top-1 predictions, except for the SSD model, which showed a $2.9\%$ difference in top-1 prediction. Although top-1 accuracy is often prioritized, it is important to recognize that changes in lower-ranked predictions can also indicate issues. Our methodology identified these inconsistencies, which have been reported to the ONNX Optimizer developers.
Figure 2: Illustrative example of problematic label classification and object detection between original unoptimized (blue) and optimized (orange) models. The bounding box location is affected for both bird instances, while the object label prediction is affected by model optimization on just the right bird instance, changing the label from bird to jet.
TABLE II: Output label accuracy differences across classification (CL) and object detection (OD) models, showing the type, opset versions, and accuracy differences for top-1, top-5, and top-10 output predictions, between the original and optimized (using the default ONNX Optimizer setting) versions of each model, along with the issue related to the observed differences.
# Model Type Opset(s) Top 1 Top 5 Top 10 Issue
1 AlexNet [30] CL 7-9 0% 0% 0.02% Format Version Change
2 DenseNet121 [25] CL 7-9 0% 0% 0.02% Format Version Change
3 GoogleNet [62] CL 7-9 0% 0% 0.04% Format Version Change
4 Inception-1 [61] CL 7-9 0% 0% 0.25% Format Version Change
5 Inception-2 [61] CL 7-9 0% 0.02% 0.05% Format Version Change
6 MobileNet v2 [24, 56] CL 7-9 0% 0% 0.02% Format Version Change
7 ResNet18 [22] CL 7 0% 0.02% 0.02% Format Version Change
8 ResNet34-v2 [23] CL 7 0% 0.02% 0.02% Format Version Change
9 ResNet50-caffe2 [22] CL 7-9 0% 2% 13% Format Version Change
10 ShuffleNet-v1 [74] CL 7-9 0% 0.7% 1% Format Version Change
11 SqueezeNet 1.0 [26] CL 7-9 0% 0% 0.02% Format Version Change
12 EfficientNet-Lite4 [63] CL 11 0% 0.04% 1% Graph Malformation
13 SSD [39] OD 12 2.9% 4.3% 6% Graph Malformation
Overall, we detected two primary issues related to output label differences between the original and optimized models:
Implicit ONNX format version upgrade:
@@ 9 ResNet50-caffe2 [22] CL 7-9 0% 2% 13% Format Version Change
The ONNX Optimizer silently upgrades models from ONNX format version v3 to v4, without any warning or indication. This behavior impacts models using opset versions 7, 8, and 9— that are older opsets but still supported in the current ONNX Runtime (v1.21.1). Despite being outdated, these opset versions are still prevalent, representing approximately 41% of the models in the ONNX Model Hub. While not all models within this opset range are affected, we observed a significant number that exhibited differences between their original and optimized versions. We have reported this issue to the ONNX Optimizer developers, who have acknowledged it and are investigating further.
Problematic fusion pass:
The pass fuse_bn_to_conv fuses Batch Normalization and Convolution operators within the model. In particular, the pass attempts to fuse BatchNormalization nodes (which do not update their mean or variance during inference) with their preceding Conv nodes by adjusting the weights and biases of the Conv node according to the properties of the BatchNormalization node. We observed that this pass was problematic in a number of cases, resulting in both a crashing bug and an issue where redundant nodes are introduced into the model. In the latter scenario, instead of properly merging the nodes resulting in a Conv node with adjusted weights and biases, the pass generates multiple nodes that represent the Batch Normalization formula. This behavior is illustrated on the right side of Figure 3. This formula representation is problematic as several models produced different output labels after optimization using this pass compared to their original versions. Although another user previously reported a similar issue, no details on the reproduction steps or affected models were provided. We submitted a reproducible issue to the ONNX Optimizer developers, including all relevant information and behavior details.
Table II shows label differences in top-K outputs for classification and object detection models, highlighting the causes. As an example, Figure 2 illustrates how an optimized model (orange) can misclassify and alter bounding boxes compared to the original model (blue).
Figure 3: Incorrect model graph transformation caused by the fuse_bn_to_conv pass in EfficientNet-Lite-4. The nodes affected are highlighted. The result should produce a Conv node with updated weights and biases, but the pass generates multiple nodes that represent the BatchNormalization formula.
TABLE III: Object detection metric differences across the original and optimized versions of object detection models, using the default ONNX Optimizer setting. The table presets the name, opset, average F1 score, precision, recall, mean Average Precision (MAP) and average Intersection over Union (IoU scores), for different IoU thresholds ($50$%, $75$%, and $90$%).
@@
(a) IoU threshold of 50%.
@@
(b) IoU threshold of 75%.
@@
(c) IoU threshold of 90%.
VI-C Bounding Box Differences in Object Detection Models
We observed notable differences in bounding box outputs between the original and optimized versions of object detection models. The results are presented in detail in Table III with the following main findings:
YOLOv3 Models:
Both YOLOv3 models (op.set $10$ and $12$) achieve high average IoU, F1 score, precision, and recall—all above $98$%. Although the optimized model occasionally misplaces bounding boxes, these errors do not cause misclassifications, as none were found in the top-10 label predictions.
YOLOv2 Model:
YOLOv2 shows lower IoU values, likely because it uses multiple bounding boxes per detected object. Small differences between the original and optimized models accumulate, lowering IoU. However, high F1, precision, and recall indicate that detection performance remains largely unaffected, as demonstrated by the bounding boxes corresponding to the bird on the left in Figure 2.
SSD Model:
In contrast, SSD exhibits a $4.8$% drop in F1, precision, and recall between original and optimized models, resulting in 2.92.92.92.96666% misclassification rates despite maintaining a high IoU ($>97$%). This behavior is illustrated in Figure 2 by the bounding boxes associated with the bird on the right.
All models show a substantial mAP difference of up to $14.31$% between original and optimized model versions. This suggests the optimizer may cause certain classes to be missed, affecting both object detection and semantic segmentation. Consequently, even if detected objects are classified correctly, some objects in the scene might be overlooked, which could significantly impact real-world applications.
VI-D Optimizer Warnings
We observed two issues related to the addition of redundant elements, specifically unused initializer entries introduced in the optimized models, associated with the eliminate_shape_gather and eliminate_nop_reshape passes. When executed with ONNX Runtime, these entries were ignored, triggering a warning in the format: Removing initializer ‘INITIALIZER_ID’. It is not used by any node and should be removed from the model. Upon manual inspection, we found no significant differences in model size. However, this issue warrants further investigation for potential side effects. We have reported it to the optimizer developers for additional review.